OpenStudy (anonymous):

f(x)= x-7/x+2 g(x)=-2x-7/x-3 find f(g(x)) and g(f(x))

OpenStudy (anonymous):

G(x)= -2x-7/x-1 (correction)

OpenStudy (anonymous):

prepare to do a raft of algebra

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[f(x)=\frac{x-7}{x+2}\\ g(x)=\frac{-2x-7}{x-1}\]

OpenStudy (anonymous):

that right?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ok lets to \[f(g(x))\] first

OpenStudy (anonymous):

what i am going to do is use cut and paste to put \[\frac{-2x-7}{x-1}\] everywhere i see an \(x\) in \[\frac{x-7}{x+2}\]

OpenStudy (anonymous):

\[\large f(g(x))=f(\frac{-2x-7}{x-1})=\frac{\frac{-2x-7}{x-1}-7}{\frac{-2x-7}{x-1}+2}\]

OpenStudy (anonymous):

clear what i did? that is the first step, bunch of algebra comes next

OpenStudy (anonymous):

Yes I understand so far :)

OpenStudy (anonymous):

ok now before we do the algebra, i want to let you know that a miracle will occur the reason you were given this problem is that \(f\) is the inverse of \(g\) and vice versa that means when we get done we will find \[f(g(x))=x\]

OpenStudy (anonymous):

Ok :D

OpenStudy (anonymous):

staring here \[\frac{\frac{-2x-7}{x-1}-7}{\frac{-2x-7}{x-1}+2}\] get rid of the compound fraction by multiplying top and bottom by \(x-1\) (carefully)

OpenStudy (anonymous):

\[\frac{-2x-7-7(x-1)}{-2x-7+2(x-1)}\] is step one

OpenStudy (anonymous):

notice the judicious use of parentheses

OpenStudy (anonymous):

now remove the parentheses using the distributive property to get \[\frac{-2x-7-7x+7}{-2x-7+2x-2}\]

OpenStudy (anonymous):

combine like terms, and since they are inverses you will get an orgy of cancellation \[\frac{-9x}{-9}=x\]finished

OpenStudy (anonymous):

method for doing the second one is similar

OpenStudy (anonymous):

That makes sense

OpenStudy (anonymous):

I am a bit confused for the second one

OpenStudy (anonymous):

same as the first, only inside out

OpenStudy (anonymous):

I put..g(f(x))= -2(x-7/x+2)-7/(x-7/x+2)-1

OpenStudy (anonymous):

try it and see what happens it will be easy enough to know if you are right or not, because the answer will be \(x\)

OpenStudy (anonymous):

yeah that looks right

OpenStudy (anonymous):

However after this step I wasn't positive if I should multiply both by x+2 or distribute

OpenStudy (anonymous):

yes multiply top and bottom by \(x+2\) like i did before, carefully, using parentheses

OpenStudy (anonymous):

Ok I'll try that :)

OpenStudy (anonymous):

I don't think I'm doing this correctly ::

OpenStudy (anonymous):

:/